OpenStudy (anonymous):

Multiply and simplify (4) (-3/2)

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

4 * (-3/2) =2 *(-3)= -6